OpenStudy (anonymous):

Maria has three children. There is two years age difference between each child. The total ages of all three children is 36 years. Rosa is the youngest child. How old is rosa? PLEASE NO CONFUSING MATH QUESTIONS JUST PLEASE HELP ME OR DONT BOTHER WRITING I NEED THIS DONE!

OpenStudy (anonymous):

Rosa is 10. -- further-- This is calculated by working out the relationship between the children. Let's call Rosa "R" and let that stand for her age. The next child is R + 2 years old and the eldest if R + 4. We are told that R + (R + 2) + (R + 4) = 36 (add the Rs and the digits on the left) R + R + R + 2 + 4 = 36 3R + 6 = 36 (subtract 6 from both sides) 3R = 36 - 6 3R = 30 (divide both sides by 3) R = 10